Calculate the bond length of C-X bond if C-C bond length is 1.54 nm and X-X bond length is 1.2 nmand electronegativities of C and X are 2.0 and 3.0 respectively.

Grade:12th pass

1 Answers

Chhavi Jain
6 years ago
Dear student, 
 
We will use the formula :
 
dA-B =  rA +  r –  0.09 | xA – xB |
where xand xB  are the values of electronegativities of A and B.
 
Now, if  dC-C = 1.54 nm  ; so rC = 1.54 / 2 = 0.77 nm
Similarly,  dx-x = 1.2nm ; so r = 1.2/ 2 = 0.6 
 
keeping the values in the above formula:
 
dc-x = 0.77 + 0.6 –  0.09 | 2-3 |
 
dc-x  = 1.28 nm   ( bond length of C-X bond )
 
Chemistry Consultant
Askiitians